OpenStudy (anonymous):

Jeff is designing the seating arrangement for a concert in his high school theater. To give everyone a good view, each row must have 5 more seats than the row before it, and the first row can only have 12 seats. Help Jeff plan the rest of the seating by solving for how many seats are in row 25. Then explain to Jeff how to create an equation to predict the number of seats in any row. Show your work, and use complete sentences. @IMStuck @aum @TinkerbellGirl @Preetha

OpenStudy (anonymous):

@hero @Australopithecus @SithsAndGiggles @Asevilla5

OpenStudy (anonymous):

Please explain?

OpenStudy (aum):

It is an arithmetic sequence with first term 12 and common difference of 5.

OpenStudy (asevilla5):

Row one= 12 Row two= 12 + 5 The equation will be: Row n = 12+5(n-1) Where n is random number, the row will have 12 seats plus 5 for every row higher than 1. So row 25 will have 12 + 5*(25-1) 12+24*5=12+120=132 Row 25 will have 132 seats. Hope it helped.

OpenStudy (anonymous):

That was quick @Asevilla5 no wonder why you call yourself the human calculator

OpenStudy (asevilla5):

I tend to be good at word problems @SCARLET_SPIDER it find it weird lol